How do you solve the following equation using the Elimination Method:

-5x + 12y = 20
x - 2y = -6

Answer by robertb(5830)

Multiply the bottom equation by 5, with the purpose of eliminating x. Then we get 5x -10y = -30. Add corresponding sides with the top equation, to get
12y - 10y = 20 - 30, or 2y = -10, or y = -5. Use the bottom equation to solve for x:
x - 2(-5) = -6, or x + 10 = -6, or x = -16. Therefore the solution is the ordered pair (-16, -5).